OpenStudy (anonymous):

Write a proportion that can be used to find the cost of 10 notebooks if 3 notebooks cost $1.98. (2 points)

hero (hero):

3/1.98 = 10/x 3x = 19.8 x = 19.8/3 x = 6.6 $6.60

OpenStudy (anonymous):

Answer: 3/1.98 = 10/x; $6.60

OpenStudy (anonymous):

Whateva

OpenStudy (anonymous):

I swear I didn't copy you. Snitch!